# Accurate Segmentation of Optic Disc and Cup from Multiple Pseudo-labels by Noise-Aware Learning

### 30 Nov 2023 v1

@@ Abstract
Optic disc and cup segmentation play a crucial role in automating the screening and diagnosis of optic glaucoma. While data-driven convolutional neural networks (CNNs) show promise in this area, the inherent ambiguity of segmenting object and background boundaries in the task of optic disc and cup segmentation leads to noisy annotations that impact model performance. To address this, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation. Specifically, the Multiple Pseudo-labels Generation and Guided Denoising (MPGGD) module generates pseudo-labels by multiple different initialization networks trained on true labels, and the pixel-level consensus information extracted from these pseudo-labels guides to differentiate clean pixels from noisy pixels. The training framework of the MPNN is constructed by a teacher-student architecture to learn segmentation from clean pixels and noisy pixels. Particularly, such a framework adeptly leverages (i) reliable and fundamental insights from clean pixels and (ii) the supplementary knowledge within noisy pixels via multiple perturbation-based unsupervised consistency. Compared to other label-denoising methods, comprehensive experimental results on the RIGA dataset demonstrate our method’s excellent performance and significant denoising ability. The code is available at https://github.com/wwwtttjjj/MPNN
@@ I Introduction
Optic disc and cup segmentation is an important step in the automated screening and diagnosis of optic nerve head abnormalities such as glaucoma. With the rapid development of deep learning, various models based on convolutional neural networks (CNNs) have emerged, showcasing promising performance in medical image segmentation. These methods [1, 2, 3, 4, 5] leverage their inherent capacity to learn intricate patterns and features from extensive labeled data, enabling them to achieve high levels of accuracy in segmenting medical images. However, due to the inherent ambiguity of the true segmentation boundaries in optic disc and cup segmentation [6, 7] and the inevitable errors in manual labeling, the annotation often contains noise. Numerous studies [8, 9, 10, 11] have indicated that label noise can significantly impact the accuracy of models.
@@
In order to mitigate the impact of label noise on medical image segmentation, a strategy is to use multiple expert annotations. Some progress has been achieved in the joint annotation learning from multiple experts for the task of optic disc and optic cup segmentation. Ji et al. [12] introduced MRNet, an innovative technique that capitalizes on the specialized insights of individual raters, treating their expertise as prior knowledge to extract high-level semantic features. Their proposed approach involves reconstructing multi-rater rankings based on initial predictions, subsequently harnessing the varying degrees of agreement among multiple raters to enhance the overall segmentation performance. By amalgamating expert knowledge and effectively pooling input from multiple raters, MRNet presents a promising avenue for improving the precision of optic disc and cup segmentation. Self-Calib [13] attempts to find ground truths by running divergent and convergent models in loops. The latter category of methods demonstrates a greater capacity for exploiting information from multiple annotations, resulting in a more impartial segmentation model.
However, it is worth noting that obtaining segmentation annotations from experts is a challenging task, not to mention the need for annotations from multiple experts. Therefore, it is necessary to explore other available methods.
@@
Without introducing additional expert labels, a series of label-denoising strategies have been proposed to cope with the challenges posed by noisy labels in segmentation tasks. These strategies typically involve two steps: first, noisy labels are identified using various methods, and then pseudo-labels are used to refine these noisy labels or to reject them. Some methods [8, 14, 15] use Confident Learning [16] modules to find noisy labels and use refinement module for processing. The method based on Confidence Learning assesses the classification confidence of each pixel’s output from the network and identifies potential noisy pixels based on this confidence level. Subsequently, the refinement module uses pseudo-labels to replace these noisy labels. Other methods [17, 18, 19, 20, 21, 22, 23] involve training multiple networks and selecting labels with the minimum loss consensus between network predictions for training. In essence, these selected pixels are treated as clean pixels, thereby eliminating noisy pixels. Compared to multi-annotation learning, the direct denoising approach is simple, efficient, and does not require the introduction of additional expert knowledge. However, while these methods have shown promise in various domains, their adaptation and implementation for the specific task of optic disc and cup segmentation still require careful investigation.
@@
In this paper, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation. To address the issue of label noise, we propose the Multiple Pseudo-labels Generation and Guided Denoising (MPGGD) module. We use multiple networks with different initialization to fit the same training set, and after reaching a certain threshold stop training to obtain multiple pseudo-labels for the training set. Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered as clean pixels and the rest are considered as noisy pixels. Furthermore, the proposed MPNN consists of a teacher-student architecture, where the student network learns by (i) minimizing the segmentation loss on the set of clean pixels and (ii) minimizing the consistency loss on the set of noisy pixels in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA). Our contributions are as follows:
@@
    To the best of our knowledge, the proposed MPNN is the first method to improve the accuracy of the optic disc and cup from the perspective of label denoising.
@@
    We proposed the MPGGD as the first execution module. It separates reliable and unreliable information, provides accurate priors for the segmentation network, and avoids the negative impact of label noise on the network.
Figure 1: Illustration of our method. The upper part is the overall network architecture, and the lower part is the MPGGD module. The MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its M_DSC on true labels $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{c}$ from the noisy pixels set $\mathbf{O}_{n}$ by consistent information.The MPNN is constructed by a model consisting of a student network and a teacher network. Where the student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{c}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{n}$ in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA).
@@ II Methods
To simplify the description of our methodology, we define $\mathbf{D}=\{\mathbf{X}_{(i)},\mathbf{Y}_{(i)}\}_{i=1}^{N}$ to represent the dataset, where $N$ represent the total samples. $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}$ denotes that the image $\mathbf{X}$ contains $n=w\times h$ pixels and $\mathbf{Y}=\{y_{1},y_{2},...,y_{n}\}$ indicates the corresponding label. Here $y\in\{0,1,2\}$ and the categories in $\{0,1,2\}$ denote the background, optic disc, and optic cup respectively. Fig. 1 illustrates our method. Specifically, the MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its M_DSC (Mean DSC of optic disc and cup) on true labels $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{c}$ from the noisy pixels set $\mathbf{O}_{n}$ by consistent information. The MPNN is constructed by a model consisting of a student network and a teacher network. Where the student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{c}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{n}$ in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA).
@@ II-A Multiple Pseudo-labels Generation and Guided Denoising (MPGGD)
Deep learning models can memorize easy samples and then gradually adapt to more difficult instances as the number of training epochs increases. When noisy pixels are present, the deep learning model ends up remembering these incorrect labels, leading to poor generalization performance. Moreover, this phenomenon does not change with training optimization or network architecture choices. Therefore, We contemplate halting the training of the network before it fits the noise and generating pseudo-labels corresponding to the training set. These pseudo-labels are then utilized to distinguish between clean and noisy labels.
@@
In order to distinguish clean pixels from noisy pixels, we use multiple networks with different initializations to fit the same training set, and after M_DSC reaches a certain threshold $\varphi$ stop training to obtain multiple pseudo-labels for the training set. Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered as clean pixels and the rest are considered as noisy pixels.
@@
Formally, give an image $\mathbf{X}\in\mathbf{D}$, the ground truths are represented $\mathbf{Y}$. We construct a set of $K$ networks $\{LinkNet(1),...,LinkNet(K)\}$ to generate multiple pseudo-labels $\mathbf{Y}_{p}=\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$. All networks are initialized with different values, taking the average dice of the optic disc and cup (M_DSC) as metric. Here $K$ and $\varphi$ are set to 5 and 0.93.
@@
In order to differentiate between noisy pixels and clean pixels, we follow the following process:
$\mathbf{O}_{c}=\{i\mid\mathbf{Y}_{p1}^{i}=\mathbf{Y}_{p2}^{i}=...=\mathbf{Y}_{pK}^{i}\}_{i=0}^{n}.$ (1)
@@
Here, $\mathbf{O}_{c}$ denotes the separated clean pixels set, and the total number of pixels contained in the set of clean pixels is denoted as $s_{c}=|\mathbf{O}_{c}|$. Noisy pixels are those in the entire set excluding clean pixels, and the collection of noisy pixels can be represented as follows:
$\mathbf{O}_{n}=\{i\mid i\notin\ \mathbf{O}_{c}\}_{i=0}^{n}.$ (2)
@@
The total number of pixels contained in the noisy pixels set is denoted as $s_{n}=|\mathbf{O}_{n}|$. These pieces of information will be used to train our network.
@@
TABLE I: Quantitative Results for Different Strategies on the RIGA Test Set. The Test Set Is Set as Average Weight Majority Vote And Individual Rater1
@@
II-B Pixel-level Semi-supervised Pattern Learning
Inherent ambiguity in segmenting objects and background boundaries in the optic disc and cup segmentation task leads to noisy annotations that impact model performance. Therefore, we study the segmentation task of the optic disc and cup with noisy labels and our method draws on the Mean-Teacher [25] model from the classical semi-supervised approach. The student network is parameterized by $\theta$ (updated by back-propagation) and the teacher network is parameterized by $\widetilde{\theta}$ (updated according to the weights $\theta$ at different training stages). The specific computation is the EMA. At training step $t$, $\widetilde{\theta}$ is updated as $\widetilde{\theta}_{t}=\alpha\widetilde{\theta}_{t-1}+(1-\alpha)\theta_{t}$, where $\alpha$ is the EMA decay rate, and it is set to 0.99, as recommended by [25].
$\mathbf{P}_{t}=\frac{1}{M}\sum_{m=1}^{M}{\mathbf{P}_{m}}.$ (3)
II-B1 Clean Pixels Loss
clean pixels provide the network with reliable and accurate training signals, thus supporting the networks’ ability to learn accurate feature representations and generalization. Our approach extracts a set of clean pixels, enabling the separation of clean and noisy information. Specifically,the clean pixels loss $\mathcal{L}_{c}$ expressed as:
$\mathcal{L}_{c}=\frac{1}{s_{c}}\sum_{i=0}^{s_{c}}{\ell_{ce}(\mathbf{P_{s}}^{\mathbf{O}_{c}^{i}},\mathbf{Y}^{\mathbf{O}_{c}^{i}}}).$ (5)
@@ II-B2 Noisy Pixels Loss
We choose entropy as a metric for estimating uncertainty. When a pixel label tends to be clean, its predictive probability distribution is likely to be peak, indicating low entropy and low uncertainty. Conversely, if a pixel label tends to be noisy, it is likely to have a flatter probability distribution, indicating high entropy and high uncertainty. Therefore, we consider the uncertainty of each pixel as a pixel noise estimate. The set of noisy pixels we separated always contains some easily predictable pixels, and the use of the uncertainty method is to find those clean pixels that can be mistaken as noisy pixels. The consistency loss in the set of noisy pixels is:
$\mathcal{L}_{n}=\frac{\sum_{i=0}^{s_{n}}\mathbbm{I}({u^{\mathbf{O}_{n}^{i}}<H})||\mathbf{P}_{s}^{\mathbf{O}_{n}^{i}}-\mathbf{P}_{t}^{\mathbf{O}_{n}^{i}}||^{2}}{{\textstyle\sum_{i=0}^{s_{n}}}\mathbbm{I}({u^{\mathbf{O}_{n}^{i}}<H})},$ (6)
where $\mathbbm{I}$ is the indicator function and $H$ (initial value is 0.75 and increases with the number of iterations) is a threshold to select the most certain targets. $u^{\mathbf{O}_{n}^{i}}$ is the estimated uncertainty $U$ at $\mathbf{O}_{n}^{i}$ pixel. Note that we only perform the selection of pixels with thresholds higher than $H$ in the noisy pixels set $\mathbf{O}_{n}$.

### 15 Mar 2024 v2

@@ Abstract
Optic disc and cup segmentation plays a crucial role in automating the screening and diagnosis of optic glaucoma. While data-driven convolutional neural networks (CNNs) show promise in this area, the inherent ambiguity of segmenting objects and background boundaries in the task of optic disc and cup segmentation leads to noisy annotations that impact model performance. To address this, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation. Specifically, the Multiple Pseudo-labels Generation and Guided Denoising (MPGGD) module generates pseudo-labels by multiple different initialization networks trained on true labels, and the pixel-level consensus information extracted from these pseudo-labels guides to differentiate clean pixels from noisy pixels. The training framework of the MPNN is constructed by a teacher-student architecture to learn segmentation from clean pixels and noisy pixels. Particularly, such a framework adeptly leverages (i) reliable and fundamental insight from clean pixels and (ii) the supplementary knowledge within noisy pixels via multiple perturbation-based unsupervised consistency. Compared to other label-denoising methods, comprehensive experimental results on the RIGA dataset demonstrate our method’s excellent performance. The code is available at https://github.com/wwwtttjjj/MPNN.
@@ I Introduction
Optic disc and cup segmentation is an important step in the automated screening and diagnosis of optic nerve head abnormalities such as glaucoma. With the rapid development of deep learning, various models based on convolutional neural networks (CNNs) have emerged, showcasing promising performance in medical image segmentation. These methods [1, 2, 3] leverage their inherent capacity to learn intricate patterns and features from extensive labeled data, enabling them to achieve high levels of accuracy in segmenting medical images. However, due to the inherent ambiguity of the true segmentation boundary in optic disc and cup segmentation and the inevitable errors in manual labeling, the annotation often contains noise. Numerous studies [4, 5] have indicated that label noise can significantly impact the accuracy of models.
@@
To mitigate the impact of label noise on medical image segmentation, one strategy is to use multiple expert annotations. Ji et al. [6] introduced MRNet, an innovative technique that capitalizes on the specialized insight of individual raters, treating their expertise as prior knowledge to extract high-level semantic features. Self-Calib [7] attempts to find ground truths by running divergent and convergent models in loops, demonstrating a greater capacity for exploiting information from multiple annotations, which results in a more impartial segmentation model. However, it is worth noting that obtaining segmentation annotations from multiple experts is a challenging task. Therefore, it is necessary to explore other available methods.
@@
Without introducing additional expert labels, a series of label-denoising strategies have been proposed to cope with the challenges posed by noisy labels in segmentation tasks. These strategies typically involve two steps: first, noisy labels are identified using various methods, and then there is the option of discarding them or refining them using pseudo-labels. Some methods [8, 9, 10, 11] involve training multiple networks and selecting labels with the minimum loss consensus between network predictions for training. Other methods [4, 12, 13] use Confident Learning [14] modules to find noisy labels and use refinement module for processing. However, their research often involves artificially simulated noisy labels and annotations from non-experts. There has not been specific research on annotation noise caused by ambiguous boundaries in the segmentation of optic disc and cup.
@@
In this paper, we propose an innovative label-denoising method of Multiple Pseudo-labels Noise-aware Network (MPNN) for accurate optic disc and cup segmentation. To address the issue of label noise, we propose the Multiple Pseudo-labels Generation and Guided Denoising (MPGGD) module. We use multiple networks with different initialization to fit the same training set, and after reaching a certain threshold stop training to obtain multiple pseudo-labels for the training set. Subsequently, the pixels corresponding to the labels that perform consistently on all the pseudo-labels are considered clean pixels, and the rest are considered noisy pixels. Furthermore, the proposed MPNN consists of a teacher-student architecture, where the student network learns by (i) minimizing the segmentation loss on the set of clean pixels and (ii) minimizing the consistency loss on the set of noisy pixels in the teacher model with multiple uncertainties. The teacher network is updated based on the parameters of the student network using the Exponential Moving Average (EMA). Our contributions are as follows:
@@
    To the best of our knowledge, the proposed MPNN is the first method to improve the segmentation accuracy of the optic disc and cup from the perspective of label denoising.
@@
    We propose the MPGGD module, which separates reliable and unreliable information, provides accurate prior information for the network and avoids the negative impact of label noise on the network.
Figure 1: Illustration of our method. (a) the architecture of MPNN; (b) the MPGGD module. The MPGGD generates multiple pseudo-labels $\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$ by multiple different initialization networks ($\{LinkNet(1),...,LinkNet(K)\}$) to fit its $DSC_{m}$ on true label $\mathbf{Y}$ to a certain threshold $\varphi$, which subsequently distinguishes the clean pixels set $\mathbf{O}_{cl}$ from the noisy pixels set $\mathbf{O}_{no}$ by consistent information. The MPNN is constructed by a model consisting of a student network and a teacher network. The student network learns by (i) minimizing the segmentation loss guided by $\mathbf{O}_{cl}$ and (ii) minimizing the consistency loss guided by $\mathbf{O}_{no}$ in the teacher model with multiple uncertainties. The teacher network is updated using the EMA.
@@ II Methods
To simplify the description of our methodology, we define $\mathbf{D}=\{\mathbf{X}_{(i)},\mathbf{Y}_{(i)}\}_{i=1}^{N}$ to represent the dataset, where $N$ represent the total samples. $\mathbf{X}=\{x_{1},x_{2},...,x_{n}\}$ denotes that the image $\mathbf{X}$ contains $n=w\times h$ pixels and $\mathbf{Y}=\{y_{1},y_{2},...,y_{n}\}$ indicates the corresponding labels. Here $y\in\{0,1,2\}$ and the categories in $\{0,1,2\}$ denote the background, optic disc, and optic cup respectively. Fig. 1 illustrates the method of MPNN. More details will be described in subsequent sections.
@@ II-A Multiple Pseudo-labels Generation and Guided Denoising (MPGGD)
Deep learning models can memorize easy samples and then gradually adapt to more difficult instances as the number of training epochs increases. When noisy labels are present, the deep learning model ends up remembering these incorrect labels, leading to poor generalization performance. Therefore, we contemplate halting the training of the network before it fits the noise and generating multiple pseudo-labels corresponding to the training set. These pseudo-labels are utilized to distinguish between clean and noisy pixels.
@@
Specifically, to distinguish clean pixels from noisy pixels, we use multiple networks with different initializations to fit the same training set, and after $DSC_{m}$ (Mean Dice Similarity Coefficient of the optic disc and cup) reaches a certain threshold $\varphi$ stop training to obtain multiple pseudo-labels of the training set. To maximize the identification of annotation noise, we adopt an over-correction strategy, that is, pixels corresponding to labels that behave consistently across all pseudo-labels are considered clean pixels, and the rest are considered noisy pixels.
@@
Formally, give an image $\mathbf{X}\in\mathbf{D}$, the ground truth is represented $\mathbf{Y}$. We construct a set of $K$ networks $\{LinkNet(1),...,LinkNet(K)\}$ to generate multiple pseudo-labels $\mathbf{Y}_{p}=\{\mathbf{Y}_{p1},...,\mathbf{Y}_{pK}\}$. All networks are initialized with different values, taking the $DSC_{m}$ as the metric. Considering the scarcity of annotation noise, here $K$ and $\varphi$ are set to 5 and 0.93.
@@
To differentiate between clean pixels and noisy pixels, we follow the following process:
$\mathbf{O}_{cl}=\{i\mid\mathbf{Y}_{p1}^{i}=\mathbf{Y}_{p2}^{i}=...=\mathbf{Y}_{pK}^{i}\}_{i=0}^{n}.$ (1)
@@
Here, $\mathbf{O}_{cl}$ denotes the separated clean pixels set, and the total number of pixels contained in the set of clean pixels is denoted as $s_{cl}=|\mathbf{O}_{cl}|$. Obviously, the noisy pixels set can be represented as follows:
$\mathbf{O}_{no}=\{i\mid i\notin\ \mathbf{O}_{cl}\}_{i=0}^{n}.$ (2)
@@
The total number of pixels contained in the noisy pixels set is denoted as $s_{no}=|\mathbf{O}_{no}|$. These pieces of information will be used to guide the training of the MPNN.
II-B Pixel-level Semi-supervised Pattern Learning
Inherent ambiguity in segmenting objects and background boundaries in the optic disc and cup segmentation task leads to noisy annotations that impact model performance. Therefore, we study the segmentation task of optic disc and cup with noisy labels and our method draws on the Mean-Teacher [15] model from the classical semi-supervised approach. The student network is parameterized by $\theta$ (updated by back-propagation) and the teacher network is parameterized by $\widetilde{\theta}$ (updated according to the weights $\theta$ at different training stages). The specific computation is the EMA. At training step $t$, $\widetilde{\theta}$ is updated as $\widetilde{\theta}_{t}=\alpha\widetilde{\theta}_{t-1}+(1-\alpha)\theta_{t}$, where $\alpha$ is the EMA decay rate, and it is set to 0.99, as recommended by [15].
$\mathbf{P}_{t}=\frac{1}{M}\sum_{m=1}^{M}{\mathbf{P}_{m}}.$ (3)
II-B1 Clean Pixels Loss
Clean pixels provide the network with reliable and accurate training signals, thus supporting the networks’ ability to learn accurate feature representations and generalization. Our approach extracts a set of clean pixels, enabling the separation of clean and noisy information. Specifically, the clean pixels loss $\mathcal{L}_{cl}$ expressed as:
$\mathcal{L}_{cl}=\frac{1}{s_{cl}}\sum_{i=0}^{s_{cl}}{\ell_{ce}(\mathbf{P_{s}}^{\mathbf{O}_{cl}^{i}},\mathbf{Y}^{\mathbf{O}_{cl}^{i}}}).$ (5)
@@ II-B1 Clean Pixels Loss
TABLE I: Comparative Results of Different Strategies on the Majority Vote Test Set and the Rater1 Test Set, Respectively
@@ II-B2 Noisy Pixels Loss
We choose entropy as a metric for estimating uncertainty. When a pixel label tends to be clean, its predictive probability distribution is likely to be peak, indicating low entropy and low uncertainty. Conversely, if a pixel label tends to be noisy, it is likely to have a flatter probability distribution, indicating high entropy and high uncertainty. Therefore, we consider the uncertainty of each pixel as a pixel noise estimate. Considering the MPGGD method is an over-correction strategy, the set of noisy pixels we separated always contains some easily predictable pixels, and the use of the uncertainty method is to find those clean pixels that can be mistaken as noisy pixels. The consistency loss $\mathcal{L}_{no}$ in the set of noisy pixels expressed as:
$\mathcal{L}_{no}=\frac{\sum_{i=0}^{s_{no}}\mathbbm{I}({u^{\mathbf{O}_{no}^{i}}<H})||\mathbf{P}_{s}^{\mathbf{O}_{no}^{i}}-\mathbf{P}_{t}^{\mathbf{O}_{no}^{i}}||^{2}}{{\textstyle\sum_{i=0}^{s_{no}}}\mathbbm{I}({u^{\mathbf{O}_{no}^{i}}<H})},$ (6)
where $\mathbbm{I}$ is the indicator function and $H$ (initial value is 0.75 and increases with the number of iterations) is a threshold to select the most certain targets. $u^{\mathbf{O}_{no}^{i}}$ is the estimated uncertainty $U$ at $\mathbf{O}_{no}^{i}$ pixel. Note that we only perform the selection of pixels with thresholds higher than $H$ in the noisy pixels set $\mathbf{O}_{no}$.
Figure 2: Visualized segmentation results on the RIGA dataset of different methods. The ground truth mask and predicted mask are represented by red and green, respectively.